OpenStudy (anonymous):

Why, given Euclid's algorithm yielding the GCD of two numbers m and n in the form (a*m + b*n), taken a step further to a' and b', does (a'*m) equal the LCM of m and n? Is there a proof for this?

OpenStudy (amistre64):

What is the definition of LCM that uses the GCD as a factor?

OpenStudy (amistre64):

there are 2 integers a,b such that the GCD of two integers m,n can be written as: am + bn = GCD what does it mean to 'take it a step further'?

OpenStudy (anonymous):

Given that a and b can be determined using the division algorithm in the format a'' = a - a'*q (similarly for b), even though you can't divide by zero, you can still continue the a and b portions of the algorithm to get an a and b such that a*m + b*n = 0, and where a*m = LCM(m,n)

OpenStudy (anonymous):

I don't know why this works but it's consistent. I just want to know if there's a proof for it or counterproof against it so I can know if it's reliable or provably false.

OpenStudy (anonymous):

For example, given m = 15, n = 10: for a = 2 and b = -3 we get: 15(2) + 10(-3) = 30 - 30 = 0 Also, 15 * 2 happens to be the LCM of 15 and 10. But how can we prove generically that this would work for ALL m and n?

OpenStudy (amistre64):

the GCD of 15 and 10 is not 0 ... but im prolly reading that wrong.

OpenStudy (anonymous):

right the gcd is 30... but lemme try to illustrate this better 15 10 15 1 0 10 0 1 5 1 -1 are you familiar with finding a and b using this method? In this case, you can see when the GCD is 5, a=1 and b=-1 ... if you continue the table one more row, however, you get 0 (because 10 MOD 5 = 0) and a=-2, b=3. As such, |a*m| = 30, the LCM(m,n) What makes this witchcraft happen? It seems to work for all m and n, but I can't work out a proof for it

OpenStudy (amistre64):

lcm is 30, gcd is 5 ... im use to a different method 15 = 1*(10) + 5 10 = 2* (5) + 0 <-- got a zero so move 1 up gcd = 5 such that 15a + 10b = 5, running a reverse euclid alg ... 0 0 1 0-1(1) = -1 out of factors, started with 0 so the smaller value gets the -1 15a + 10(-1) = 5, a has to be 1 15(1) + 10(-1) = 5

OpenStudy (amistre64):

try your method for a prime, or relatively prime values and see if it works out for you.

OpenStudy (anonymous):

Yes, it is working for all values, even if I make m=n it woks, if m or n = 1, or both do, etc.

OpenStudy (amistre64):

can you post a link to the method you are using? it would help me better see the process and tell if things are working in some pattern that can be proofed

OpenStudy (anonymous):

I haven't found the method published anywhere, it was my professor's blackboard method. I unfortunately am having a difficult time explaining it better than this. Basically you build a table using m and n, then recursively fill out each row of the table using DIV, MOD, and new a and b values recursively until MOD = 0; that's how you know you've reached the GCD. That's Euclid's algorithm in a nutshell. But then AFTER you find GCD, if you continue filling out the table to the next row you get numbers that start giving you LCM of m and n and we can't figure out why. Sorry, I am not able to explain it better in chat format. :(

OpenStudy (amistre64):

so if im reading this right ... 15 10 15 15/15 = 1 10/15 = 0, assuming since its less than 1? 10 15/10 = 0 10/10 = 1 ^^^^ this is where my comprehension is breaking down explain how we fill in the first row, then lets see if i can use that to fill in the second row, and proceed to the third row.

OpenStudy (experimentx):

looks like something interesting is cookin up .. .what's happening??

OpenStudy (amistre64):

euclid is casuing trouble again :)

OpenStudy (experimentx):

lol ... (a*m + b*n) = gcd is Bezout lemma ... i recently did that on abstract algebra. freaking ugly.

OpenStudy (experimentx):

but what @steven_k is trying to tell ... I am not understanding, probably my english :|

OpenStudy (amistre64):

if i could figure out the tabling i might be able to perceive it regardless of my english :)

OpenStudy (anonymous):

Sorry, the method is quite complicated... The first rows are filled in as they are because 15(1) + 10(0) = 15, and 15(0) + 10(1) = 10. Then in the second row, the first number is 5 because 15mod10=5, the second number is 1 and the third number is -1 because for each number in row (n), it is (n-2) - q(n-1), where q = the DIV of the previous two row headers, in this case 15 and 10; so DIV = 1

OpenStudy (experimentx):

did you mean that how Euclid's algo works?

OpenStudy (experimentx):

*do

OpenStudy (anonymous):

yes, but the point being, using this method you can continue one row past the end where you find GCD and get values for a and b such that (a*m)+(b*n) = 0 and |a*m| = lcd(m,n)

OpenStudy (experimentx):

(a*m)+(b*n) = 0 is equivalent to (a*m)= (b*n) ... just ignore the negative, ... this is trivially true, just set m=b and n=a